# HS-FPN: High Frequency and Spatial Perception FPN for Tiny Object Detection

### v2

@@ A.1 Motivation of HFP
Figure 8: Single channel feature map visualization before and after applying high pass filter. $\alpha$ is set to 0.25. The first column and the second column show the visualizations of different channels, respectively.
As illustrated in Figure 8 (a), the richness of features related to tiny objects varies across different channels of the input feature map, often accompanied by some background information. However, after applying high-pass filtering, as shown in Figure 8 (b), the filtered result (the Hadamard product of the original features and the high-pass filtering mask) predominantly highlights the features of the tiny objects, while suppressing background-related information. This demonstrates the significance of performing high-pass filtering before the channel and spatial branches in the HFP. By increasing the proportion of tiny object features in the feature map, high-pass filtering enhances the accuracy of channel attention extraction.
Figure 9: Visualization of channel weight.
Figure 10: Visualization of high-pass filtering masks under different values of $\alpha$.
Figure 9 provides a visualization of the channel weights, offering insight into the role of the channel branch in HFP. The channel branch assigns varying weights to each channel based on its high-frequency response, ensuring that the weights in the input feature map are no longer uniform, but are dynamically allocated according to the tiny object features present in each channel.
Figure 10 presents visualizations of high-pass filtering masks for different values of $\alpha$. As $\alpha$ increases, the masks retain more high-frequency information. With an optimal value of $\alpha$ (e.g., $\alpha=0.25$), the background is effectively filtered out, while rich high-frequency details are preserved.
@@ A.2 Differences between SDP and ViT
In summary, ViT implements region-level attention, whereas SDP focuses on pixel-level attention. As a result, SDP is capable of learning spatial dependencies between pixels and their local surroundings, thereby enriching the spatial context information of tiny objects. However, it is important to note that the computational cost of SDP is $\frac{(hw)}{n}$ times that of ViT, and this cost increases as the patch size $hw$ becomes larger.
@@ A.4 Train Details
Figure 12 shows the visualization results on AI-TOD. The first and third rows are the results of Cascade R-CNN with FPN, the second and fourth rows are the results of Cascade R-CNN with HS-FPN. Backbones are both ResNet50. It can be observed that replacing FPN with HS-FPN significantly reduces both FN and FP predictions for the targets, indicating that HS-FPN improves the detection performance of tiny objects.

### v3

@@ A.1 Motivation of HFP
Figure 8: Visualization of high-frequency responses $F_{2}$ generated under different $\alpha$.
@@
Figure 8 presents the visualization of the high-frequency responses $F_{2}$ extracted from $C_{2}$ under different $\alpha$. It can be observed that when $\alpha$ is 0, all frequency bands can pass through filter $z$ making $F_{2}$ identical to the input feature $C_{2}$, which contains a large amount of low-frequency background information. As $\alpha$ gradually increases, the low-frequency components in $C_{2}$ are progressively filtered out by $z$, and the tiny object features in the generated high-frequency responses $F_{2}$ become increasingly prominent, with their proportion in $F_{2}$ significantly higher than in $C_{2}$. However, when $\alpha$ continues to increase beyond an optimal value, even the tiny object features in $C_{2}$ start to be filtered out by filter $z$, which manifests as the high-frequency noise becoming more pronounced in sub-figures (c-e), while the target features gradually weaken. When $\alpha$ reaches its maximum value of 1, $z$ completely blocks all frequency bands, resulting in the high-frequency response $F_{2}$ becoming an empty matrix, as shown in sub-figure (f).
It can be observed from Figure 9, due to the different parameters of convolutional kernels in each channel, the extracted features emphasize different aspects, leading to variations in the feature representation across channels in $C_{2}$. Notably, the proportion of tiny object features varies among different channels of $C_{2}$, which consequently affects their contribution to the final detection performance. The purpose of the channel branch is to identify channels in $C_{2}$ that contain more tiny object features and assign different weights to enhance their feature representation.
Figure 9: Input image and partial channel feature visualization of $C_{2}$.
Figure 10: Visualization of channel weight generated by HFP.
However, since tiny object features occupy only a small proportion of the original feature map, computing channel weights directly on the entire feature map is susceptible to interference from low-frequency homogeneous background features. To address this, the High Frequency Generator (HFG) in Figure 3 can be used to filter out most low-frequency components in $C_{2}$, resulting in the high-frequency response $F_{2}$. Since tiny object features belong to the high-frequency components, and the low-frequency components are removed in $F_{2}$, the proportion of tiny object features in $F_{2}$ is significantly increased compared to $C_{2}$. At this stage, we can more accurately identify the presence of tiny object features in each channel of $C_{2}$ based on $F_{2}$, thereby computing more precise channel weights.
The channel weights generated by HFP are visualized in Figure 10, where the channel branch adaptively assigns different weights to each channel based on its high-frequency response. This ensures that the weights in the input feature map are no longer uniform but are dynamically adjusted according to the tiny object features present in each channel.
Figure 11: Visualization of the feature at the $P_{2}$ level after HS-FPN.
Figure 12: Visualization of the feature at the $P_{2}$ level after HFP.
Figure 13: Visualization of the feature at the $P_{2}$ level after SDP.
Figures 11, 12, and 14 respectively present the visualization results of the P2 layer features before and after processing by HS-FPN, HFP alone, and SDP alone. It can be observed that each component of HS-FPN effectively expands the feature dimensions of tiny objects and enhances their feature responses.
@@ A.2 Differences between SDP and ViT
In summary, ViT implements region-level attention, whereas SDP focuses on pixel-level attention. As a result, SDP is capable of learning spatial dependencies between pixels and their local surroundings, thereby enriching the spatial context information of tiny objects. However, it is important to note that the computational cost of SDP is $(hw)/{n}$ times that of ViT, and this cost increases as the patch size $hw$ becomes larger.
@@ A.3 Datasets and Metrics
Figure 15: Target size distribution map of $DOTA_{mini10}$.
@@ A.4 Train Details
Figure 16, 17 shows the visualization results on AI-TOD and $DOTA_{mini10}$. The first and third rows are the results of Cascade R-CNN with FPN, the second and fourth rows are the results of Cascade R-CNN with HS-FPN. Backbones are both ResNet50. It can be observed that replacing FPN with HS-FPN significantly reduces both FN and FP predictions for the targets, indicating that HS-FPN improves the detection performance of tiny objects.
@@
Figure 17: Visualization results on $DOTA_{mini10}$. The first and third rows are the results of Cascade R-CNN with FPN, the second and fourth rows are the results of Cascade R-CNN with HS-FPN. The green, blue and red boxes in the figure donate true positive (TP), false positive (FP) and false negative (FN) predictions. Zooming in for more details.